OpenStudy (anonymous):

Is x = 5 a solution to -2x + 10 = 0? Write yes or no.

OpenStudy (y2o2):

yes

OpenStudy (y2o2):

-2(5)+10 = -10+10 = 0

OpenStudy (anonymous):

absolutely :D

OpenStudy (anonymous):

yes